# Simplicity bias, algorithmic probability, and the random logistic map

### 31 Dec 2023 v1

@@ 3.2. Adding noise to the strings
In Figure 2 we see the effects of adding measurement noise. For illustration purposes, we use the same values of $\mu$ and $\epsilon$ as in Figure 1, and increasing values of $\delta$. In Figure 2(a), a very small noise level of $\delta=0.01$ is used, and hence the simplicity bias is still clearly observed. In (b), a larger noise level of $\delta=0.17$ is used, and we see that the simplicity bias still remains, but the slope of decay in log probability with increasing complexity is slightly less steep as compared to (a). Finally, in (c), a larger noise level of $\delta=0.45$ is used, and the bias and simplicity bias has almost completely disappeared with only a modest slope decay in log probability. In sum, the simplicity bias gradually disappears with increasing $\delta$ via the slope gradually becoming more horizontal.
@@ 3.4. More data does not always give more confidence in predictions
In the preceding sections, we looked at how simplicity bias, which is a form of algorithmic probability, can be used to predict or bound the probability of patterns in time series from the logistic map. Another application of algorithmic probability, and in fact one of the original motivations for its development, is for inference or induction of series. Solomonoff [21, 24] was interested in questions like, given a (finite) binary series $x$ representing a historical series, what is the best prediction for the next bit in the series? More quantitatively, what is the probability that the next bit is a zero, $P(0|x)$, and what is the probability the next bit is a 1, $P(1|x)$? If the underlying process generating the series is known, e.g. it is a Bernoulli process, then standard statistical theory can give methods to estimate the probabilities of $P(0|x)$ and $P(1|x)$. In many practical situations, however, the underlying process is not known, and a central goal of the study of time series is forecasting using just the series itself. So, is there a general way to predict time series, which does not require (strong) assumptions or knowledge about the process that made the series? Solomonoff induction [40] is the method proposed by AIT, which is essentially prediction based on algorithmic probability.
Hutter [60] has also argued that algorithmic probability should be central to artificial intelligence and the theory of machine learning. The main problem with algorithmic probability, and hence Solomonoff induction, is that it is formally uncomputable [40], so cannot be applied directly to in practice (or at least not straightforwardly). Indeed, some have argued that this makes the theory almost useless [61]. However, developing the theory is useful because some aspects can shed light on real-world problems, and some of the theory can be approximated and thereby applied.
Here we will consider one rather surprising result from Solomonoff induction, which is that sometimes with more data — even if the data points all conform to one trend — we should be less confident in our ability to predict the series (see Li and Vitanyi [40] for a mention of this result, and more in Hutter [62]). We will see how this result can apply to the logistic map.
@@
Solomonoff induction (or prediction) is based on imagining that a computer has been fed a random input program, and it produces an infinite binary string output. We observe the first $n$ bits of the output as a binary string $x$, and we try to predict the probability that the next bit is a 0 or a 1. Technically this computer should be a monotone universal Turing machine, but we will not explore the details of this because they are not relevant to our example, see [26, 62, 40] for more details. We will write $M(x)$ for the probability that the infinite string begins with $x$, and write the conditional probability as
@@
For most $n$, or typical $n$, it is known that the complexity of an integer $n$ is given by $K(n)\approx\log_{2}(n)$ [40], so that
@@
A difference in predictions between the Laplacian and algorithmic probability approaches can arise when certain values of $n$ are encountered, which are ‘special’, and therefore (algorithmically) ‘simple’. Such values of $n$ are those that can be described succinctly and unambiguously. For example, $n=2^{2^{2^{2^{2}}}}\approx 2\times 10^{19728}$, meaning that if the integer $n$ is written out in full it has nearly 20,000 digits, while at the same time, we can describe it very succinctly in just a few characters. Therefore we say that $n$ is simple, and the Kolmogorov complexity of this value $n$ is low. As another example of ‘special’ or ‘simple’ integers, the factorial $n=(8!)!\approx 3\times 10^{168186}$, which has nearly 170,000 digits if written out explicitly in full, yet can be described exactly in just a few characters. There are many ‘special’ and algorithmically simple numbers, but nonetheless, they are rare in the sense that a vanishingly small fraction of integers are special or simple [40].
For such ‘special’ or ‘simple’ numbers $n$, the Kolmogorov complexity is small, or we can say $O(1)$ bits, to represent just a few bits. In other words, for a simple $n$ which has $K(n)\ll\log_{2}(n)$, the predicted probability for the next bit is
@@
which can be substantially different from Laplace’s estimate. This is not to say that algorithmic probability would predict a 0 in place of a 1 as the next bit, but instead, it means that the prediction that the next bit is 1 — for this $n$ — is less confident (further from 1) than it would be for a typical $n$. This is a kind of quantification of the intuition that typically we do not expect there to be a change in a series after a certain pattern has been followed for a long time (e.g., 1s to change to 0s), but on ‘special occasions’ when $n$ is somehow special, then it is more likely that something will ‘happen’ (i.e., a change).
Rather surprisingly, what this means also is that for some few special values of $n$, more data does not necessarily mean higher confidence in our predictions: consider $n$ is some large random integer such that $K(n)\approx\log_{2}(n)$, while $n^{\prime}$ is some other simple integer such that $n^{\prime}>n$ but at the same time
@@
We will now consider how this somewhat counterintuitive result from algorithmic probability theory applies to the logistic map. We will construct an example of when the trend of just 0s changes after a long time and becomes a string of 1s.
If $\mu=2.5$, $\epsilon=0.0$, $(x_{0})^{-1}=2^{2^{2^{2^{2}}}}$ so that $x_{0}\approx 10^{-19728}$, then this series begins with $x_{0}<0.5$. Because $\mu=2.5$, the fixed point which the trajectory approaches is $0.51-1/\mu=1-1/2.5=0.6>$ 0.5. Therefore, eventually, the series will change from 0s to 1s. Call $n^{*}$ the integer at which the series transitions from 0s to 1s, i.e. the series is $0000...011111$.
Because $x_{0}\approx 0$ and $\epsilon=0.0$, we have $(1-x_{k})\approx 1$ so that the logistic equation Eq. (3) simplifies to $x_{k+1}\approx 2.5x_{k}$, or $x_{k+1}\approx 2.5^{k}x_{0}$, meaning that when $x_{k}\approx 0$ the terms grow exponentially fast initially, but slows as $x_{k}$ gets larger. Even with exponential growth, $n^{*}$ must be a large number because $x_{0}$ is so small, indeed because $10^{19728}\approx 2.5^{49575}$, the number of iterations $n^{*}$ required to reach the 0.5 threshold is at least of the order of 50,000.
So, having observed a series of $n^{*}$ 0s, $x=0^{n^{*}}$, according to Laplace’s induction method, the probability that the next bit is also 0 will be around
$P(0|000...000)=\frac{k+1}{n+2}=\frac{n^{*}+1}{n^{*}+2}\approx 1-1/n^{*}>1-1/50000=0.99998$ (12)
@@
In contrast, using algorithmic probability we have the following reasoning: Because $x_{0}$ is an algorithmically simple (small) number, and $\mu$ is a simple number also, and the logistic equation Eq. (3) can be written in only a few characters also, and because the value of $n^{*}$ can be derived via a simple function of these values (e.g. by iterating the logistic map to find the value of $n^{*}$), then this implies that $n^{*}$ is a special number and $K(n^{*})$ is small. Therefore we have
$M(0|000...000)=1-2^{-K(n^{*})+O(1)}<1-1/n^{*}$ (13)
so that $M(1|000...000)>1/n^{*}$, so that the change from 0s to 1s is less of a ‘surprise’ if using the algorithmic probability approach.
@@ 4. Discussion
Arguments inspired by algorithmic information theory (AIT) have been used to predict the occurrence of simplicity bias in many real-world input-output maps, where complex patterns have exponentially low probabilities, and high probability patterns are simple [31]. This phenomenon has been observed in a very wide range of systems, including RNA shapes [31, 38], protein shapes [52, 50], models of plant growth [31], ordinary differential equation solution profiles [31], finite state transducer [32], natural time series [43], deep neural networks [33], 1D dynamical systems [27], and others. In this work, we have numerically investigated the presence of simplicity bias in digitized trajectories of the random logistic map. This current work sits in the context of a wider research program of investigating the ‘universal’ presence of simplicity bias in input-output maps, which can aid in an understanding of the origins of symmetry and simplicity in nature, and also serve the practical aim of aiding a priori probability predictions, based on directly measuring the complexity of patterns (instead of, e.g. sampling).
@@
It is noteworthy that digitized logistic map trajectories have been studied in terms of Lempel-Ziv complexity by Kaspar and Schuster [55], and in this sense is related to our work. Despite this, their work is fundamentally different from ours because it is not concerned with simplicity bias or in estimating the probability of different outputs.
The theory of algorithmic probability was developed several decades ago [21, 24, 63, 25], and is important and potentially widely applicable in science and mathematics [40, 60, 64, 26, 65]. Applications of the field have been held back, partly at least, by the difficulty of applying the theory due to the incomputability of important quantities. However, it has been argued that algorithmic probability predictions should “serve as ‘gold standards’ that practitioners should aim at … [but which] have to be (crudely) approximated in practice” [62]. The simplicity bias bound, including the results here, can be seen in this light. Further, many other researchers have built on and applied the compression-learning-prediction link [66]. Indeed, very recently it has been argued that even the currently highly influential “large language models [e.g., ChatGPT] are powerful general-purpose predictors and that the compression viewpoint provides novel insights” [67].
@@ 5. Conclusion
In this study, we’ve demonstrated the presence of simplicity bias within the random logistic map, affirming its role across a range of dynamic systems. We observed that simplicity bias persists within specific parameter regimes but is gradually diminished by measurement noise, particularly within noise-induced chaos regimes. These findings not only reinforce the intricacies of dynamical systems but also emphasize the counterintuitive nature of algorithmic probability-based predictions.
@@
As we look forward, the application of algorithmic probability and related concepts remains a promising frontier for advancing prediction and analysis methods in various fields. By continuing to bridge theory with practical applications, we aim to deepen our understanding of the principles that govern simplicity and complexity in the natural world.
Acknowledgments: BH thanks Prof. Jeroen Lamb (Imperial College London) for useful discussions about random dynamical systems that inspired the work in this paper. KD thanks Muhammad Alaskandarani for useful discussions on the early parts of this work. This work has been partially supported by the Gulf University for Science and Technology, including by project code: ISG Case 9.
@@ Appendix B Comparison of Laplace and algorithmic probability for induction
In the logistic map, assume that $\mu=2.5$ and $\epsilon=0.0$. As in the main text, if we assume that $x_{0}\approx 0$, then the digitised trajectory will begin as zeros, and then change to 1s as the trajectory approaches $1-1/\mu=$ 0.6$>$0.5. Define $x_{0}$ to be the value such that the series of 0s changes to 1s after $m^{m}$ zeros are observed, where $m$ is some positive integer. Because the logistic map can be described in a few bits, and $\mu=2.5$ and $\epsilon=0.0$ can also be described in a few bits, then the complexity of the series comes from the value of $m$. We have
@@
for typical $m$. Therefore the algorithmic probability prediction that the next bit after the zeroes will be 1 is $2^{-\log_{2}(m)}\sim 1/m$. In contrast, after observing $m^{m}$ zeros, the Laplacian prediction for the probability that the next bit is 1 is $1/m^{m}$ which is substantially smaller. Therefore the observation of a 1 is more ‘surprising’ from the perspective of the Laplacian method.

### 08 Apr 2024 v2

@@ 3.2. Adding noise to the strings
In Figure 2 we see the effects of adding measurement noise. For illustration purposes, we use the same values of $\mu$ and $\epsilon$ as in Figure 1, and increasing values of $\delta$. In Figure 2(a), a very small noise level of $\delta=0.01$ is used, and hence the simplicity bias is still clearly observed. In (b), a larger noise level of $\delta=0.17$ is used, and we see that the simplicity bias still remains, but the slope of decay in log probability with increasing complexity is slightly less steep as compared to (a). Finally, in (c), a larger noise level of $\delta=0.45$ is used, and the bias and simplicity bias has almost completely disappeared with only a modest slope decay in log probability. The slopes of the upper bounds are -0.12, -0.10, and -0.04 respectively. In sum, the simplicity bias gradually disappears with increasing $\delta$ via the slope gradually becoming more horizontal.
@@ 3.4. More data does not always give more confidence in predictions
In the preceding sections, we looked at how simplicity bias, which is a form of algorithmic probability, can be used to predict or bound the probability of patterns in time series from the logistic map. Another application of algorithmic probability, and in fact one of the original motivations for its development, is for inference or induction of series. Solomonoff [23, 26] was interested in questions like, given a (finite) binary series $x$ representing a historical series, what is the best prediction for the next bit in the series? More quantitatively, what is the probability that the next bit is a zero, $P(0|x)$, and what is the probability the next bit is a 1, $P(1|x)$? If the underlying process generating the series is known, e.g., it is a Bernoulli process, then standard statistical theory can give methods to estimate the probabilities of $P(0|x)$ and $P(1|x)$. In many practical situations, however, the underlying process is not known, which motivates the development of forecasting methods which do not depend on knowing the process. So, is there a general way to predict time series, which does not require (strong) assumptions or knowledge about the process that made the series? Solomonoff induction [42] is the method proposed by AIT, which is essentially prediction based on algorithmic probability.
Hutter [62] has also argued that algorithmic probability should be central to artificial intelligence and the theory of machine learning. The main problem with algorithmic probability, and hence Solomonoff induction, is that it is formally uncomputable [42], so cannot be applied directly in practice (or at least not straightforwardly). Indeed, some have argued that this makes the theory almost useless [63]. However, developing the theory is useful because some aspects can shed light on real-world problems, and some of the theory can be approximated and thereby applied.
Here we will consider one rather surprising implication of Solomonoff induction: sometimes with more data — even if the data points all conform to one trend — we should be less confident in our ability to predict the series (see Li and Vitanyi [42] for a mention of this result, and more in Hutter [64]). We will see how this result can apply to the logistic map. To our knowledge, the logistic map example presented below is the first concrete example illustrating this implication.
3.4.1. Theory
@@
Solomonoff induction (or prediction) is based on imagining that a computer has been fed a random input program, and it produces an infinite binary string output. We observe the first $n$ bits of the output as a binary string $x$, and we try to predict the probability that the next bit is a 0 or a 1. Technically this computer should be a monotone universal Turing machine, but we will not explore the details of this because they are not essential to understanding our example; see [28, 64, 42] for more details. We will write $M(x)$ for the probability that the infinite string begins with $x$, and write the conditional probability as
@@
For most $n$, or typical $n$, it is known that the complexity of an integer $n$ is given by $K(n)\approx\log_{2}(n)$ to first order [42], so that
@@
A difference in predictions between the Laplacian and algorithmic probability approaches can arise when certain values of $n$ are encountered, which are ‘special’, and therefore (algorithmically) ‘simple’. Such values of $n$ are those that can be described succinctly and unambiguously. For example, if $n=2^{2^{2^{2^{2}}}}\approx 2\times 10^{19728}$ is written out in full it has nearly 20,000 digits (in base 10), while at the same time we can describe it very succinctly in just a few characters using a stacked exponential. Therefore we say that $n$ is simple, and the Kolmogorov complexity of this value $n$ is low. As another example of ‘special’ or ‘simple’ integers, consider the factorial $n=(8!)!\approx 3\times 10^{168186}$. This integer has nearly 170,000 digits if written out explicitly in full, yet can be described exactly in just a few characters. There are many ‘special’ and algorithmically simple numbers, but nonetheless, they are rare in the sense that a vanishingly small fraction of integers are special or simple [42].
For such ‘special’ or ‘simple’ numbers $n$, the Kolmogorov complexity is small. It follows that for a simple $n$ which has $K(n)\ll\log_{2}(n)$, the predicted probability for the next bit is
@@
which can be substantially different from Laplace’s estimate for large enough $n$. This is not to say that algorithmic probability would predict a 0 in place of a 1 as the next bit, but instead, it means that the prediction that the next bit is 1 — for this $n$ — is less confident (further from 1) than it would be for a typical $n$. This is a kind of quantification of the intuition that typically we do not expect there to be a change in a series after a certain pattern has been followed for a long time (e.g., 1s to change to 0s), but on ‘special occasions’ when $n$ is somehow special, then it is more likely that something will ‘happen’ (i.e., a change).
Rather surprisingly, what this means also is that for some few special values of $n$, more data does not necessarily mean higher confidence in our predictions: consider $n$ is some large random integer such that $K(n)\approx\log_{2}(n)$, while $n^{\prime}$ is some other simple integer such that $n^{\prime}>n$. It follows that
@@
3.4.2. Example
We will now consider how this somewhat counterintuitive result from algorithmic probability theory applies to the logistic map. We will construct an example of when the trend of just 0s changes after a long time and becomes a string of 1s at an algorithmically simple value of $n$.
Suppose we choose $\mu=2.5$, $\epsilon=0.0$, and $(x_{0})^{-1}=2^{2^{2^{2^{2}}}}$ so that $x_{0}\approx 10^{-19728}$, then this series begins with $x_{0}<0.5$. Because $\mu=2.5$, the fixed point which the trajectory approaches is $0.51-1/\mu=1-1/2.5=0.6>$ 0.5. Therefore, eventually, the digitised series will change from 0s to 1s. Call $n^{*}$ the integer at which the series $0000...011111$ transitions from 0s to 1s.
Because $x_{0}\approx 0$ and $\epsilon=0.0$, we have $(1-x_{k})\approx 1$ for $k\approx 1$ so that the logistic equation Eq. (3) simplifies to $x_{k+1}\approx 2.5x_{k}$, or $x_{k+1}\approx 2.5^{k}x_{0}$. This implies that when $x_{k}\approx 0$ the $x_{k}$ terms grow exponentially fast initially (but slows as $x_{k}$ gets larger). Even with exponential growth, $n^{*}$ must be a large number because $x_{0}$ is so small, indeed because $10^{19728}\approx 2.5^{49575}$, the number of iterations $n^{*}$ required to reach the 0.5 threshold is at least of the order of 50,000.
So, having observed a series of $n^{*}$ 0s, according to Laplace’s induction method, the probability that the next bit is also 0 will be roughly estimated as
$P(0|000...000)=\frac{k+1}{n+2}=\frac{n^{*}+1}{n^{*}+2}=1-1/n^{*}>1-1/50000=0.99998$ (12)
@@ 3.4.2. Example
In contrast, using algorithmic probability we have the following reasoning: Because $x_{0}$ is an algorithmically simple (small) number, and $\mu$ is a simple number also, and the logistic equation Eq. (3) can be written in only a few characters also, and because the value of $n^{*}$ can be derived via a simple function of these values (e.g., by iterating the logistic map to find the value of $n^{*}$), then this implies that $n^{*}$ is a special number and $K(n^{*})$ is small [^1] Technically, the value of the complexity $K(n^{*})$ will depend on the specific computer of language used, but we ignore this for this example. Later we give an example which does not depend on these.. Therefore we have
$K(n^{*})<\log_{2}(n^{*})\Rightarrow M(0|000...000)=1-2^{-K(n^{*})+O(1)}<1-1/n^{*}$ (13)
so that $M(1|000...000)>1/n^{*}$, so that the change from 0s to 1s is less of a ‘surprise’ if using the algorithmic probability approach. It follows also that we are less confident about predicting the next bit in the sequence $1^{n^{*}}=111\dots 111$, as compared to predicting the next bit in the sequence $1111\dots 1$ 111 where the length of the sequence is some algorithmically random but slightly smaller value than $n^{*}$.
@@ 4. Discussion
Arguments inspired by algorithmic information theory (AIT) have been used to predict the occurrence of simplicity bias in many real-world input-output maps, where complex patterns have exponentially low probabilities, and high probability patterns are simple [33]. This phenomenon has been observed in a very wide range of systems, including RNA shapes [33, 40], protein shapes [54, 52], models of plant growth [33], ordinary differential equation solution profiles [33], finite state transducer [34], natural time series [45], deep neural networks [35], 1D dynamical systems [29], and others. Relatedly, Zenil and Delahaye [44] numerically studied the frequencies by which very short ($\sim$5 bits) binary string motifs appear in nature, namely in hard drives, DNA fragments, and real-world image data. They found a correlation between these short binary string motifs, and the frequencies with which the same motifs appeared from generic computation devices (e.g., cellular automata). Their investigation showed that some simpler binary strings occurred with higher probability, hence a form of simplicity bias in natural data.
In this work, we have numerically investigated the presence of simplicity bias in digitized trajectories of the random logistic map. This current work sits in the context of a wider research program of investigating the ‘universal’ presence of simplicity bias in input-output maps, which can aid in an understanding of the origins of symmetry and simplicity in nature, and also serve the practical aim of aiding a priori probability predictions, based on directly measuring the complexity of patterns (instead of e.g., sampling).
@@
It is noteworthy that digitized logistic map trajectories have been studied in terms of Lempel-Ziv complexity by Kaspar and Schuster [57], and in this sense is related to our work. Despite this, their work is fundamentally different from ours because it is not concerned with simplicity bias or in estimating the probability of different outputs. Also, Zenil et al. [65] have studied discrete dynamical systems of cellular automata from the perspective of algorithmic information theory, including reconstructing the space-time evolution of the systems. However, their work is different several respects, including that they did not look at the (random) logistic map in terms of simplicity bias.
The theory of algorithmic probability was developed several decades ago [23, 26, 66, 27], and is important and potentially widely applicable in science and mathematics [42, 62, 67, 28, 68]. Applications of the field have been held back, partly at least, by the difficulty of applying the theory due to the incomputability of important quantities. However, it has been argued that algorithmic probability predictions should “serve as ‘gold standards’ that practitioners should aim at … [but which] have to be (crudely) approximated in practice” [64]. The simplicity bias bound, including the results here, can be seen in this light. Further, many other researchers have built on and applied the compression-learning-prediction link [69]. Indeed, very recently it has been argued that even the currently highly influential “large language models [e.g., ChatGPT] are powerful general-purpose predictors and that the compression viewpoint provides novel insights” [70]. It remains to be seen in the near future whether this intriguing compression approach provides valuable outcomes for machine learning.
@@ 5. Conclusion
In this study, we have demonstrated the presence of simplicity bias within the random logistic map, affirming its role across a range of dynamic systems. We observed that simplicity bias persists within specific parameter regimes but is gradually diminished by measurement noise, particularly within noise-induced chaos regimes. These findings not only reinforce the intricacies of dynamical systems but also emphasize the counterintuitive nature of algorithmic probability-based predictions.
@@
As we look forward, the application of algorithmic probability and related concepts remain a promising frontier for advancing prediction and analysis methods in various fields. By continuing to bridge theory with practical applications, we aim to deepen our understanding of the principles that govern simplicity and complexity in the natural world.
Acknowledgments: BH thanks Prof. Jeroen Lamb (Imperial College London) for useful discussions about random dynamical systems that inspired the work in this paper. BH acknowledges support from the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration and from Beyond Limits (Learning Optimal Models) through CAST (The Caltech Center for Autonomous Systems and Technologies).
KD thanks Muhammad Alaskandarani for useful discussions and work on the early parts of this work. This work has been partially supported by the Gulf University for Science and Technology, including by project code: ISG Case 9.
@@ Appendix B Comparison of Laplace and algorithmic probability for induction
In the logistic map, assume that $\mu=2.5$ and $\epsilon=0.0$. As in the main text, if we assume that $x_{0}\approx 0$, then the digitised trajectory will begin as zeros, and then change to 1s as the trajectory approaches $1-1/\mu=$ 0.6$>$0.5. Define $x_{0}$ to be the value such that the series of 0s changes to 1s after $m^{m}$ zeros are observed, where $m$ is some positive integer. Because the logistic map can be described in a few bits, and $\mu=2.5$ and $\epsilon=0.0$ can also be described in a few bits, then the largest contribution to the complexity of the series comes from the value of $m$. We have
@@
for typical and large $m$. Therefore the algorithmic probability prediction that the next bit after the zeroes will be 1 is roughly $2^{-\log_{2}(m)}\sim 1/m$. In contrast, after observing $m^{m}$ zeros, the Laplacian prediction for the probability that the next bit is 1 is $1/m^{m}$ which is substantially smaller. Therefore the observation of a 1 is more ‘surprising’ from the perspective of the Laplacian method.